Let z, w ∈ C. Show |z + w| =


p


|z|


2 + 2Re(zw) + |w|


2 =


p


|z|


2 + 2|z||w| cos(θz − θw) + |w|


2 and


|z + w| = (|z| + |w|)


s


1 −






2 sin([θz−θw]/2)


|z|/|w|+



|w|/|z|


2


, if z 6= 0 and w 6= 0.

1
Expert's answer
2021-11-10T14:20:24-0500

Let z, w ∈ C. Show

"|z + w| =\sqrt{|z|^2 + 2Re(z\overline{w}) + |w|^2} =\sqrt{|z|^2 + 2|z||w| cos(θ_z − θ_w) + |w|^2}"


if "z \neq 0 ,w \neq 0" .


"|z + w|^2=(z+w)(\overline{z+w})=(z+w)(\overline{z}+\overline{w})=z\overline{z}+w\overline{w}+z\overline{w}+\overline{z}w="

"=|z|^2+|w|^2+z\overline{w}+\overline{z}w=|z|^2+|w|^2+2Re(z\overline{w})"


"z\overline{w}=|z||w|(cos\theta_z+isin \theta_z)(cos\theta_w-isin \theta_w)"

"Re(z\overline{w})=cos\theta_z cos\theta_w+sin\theta_z sin\theta_w=|z||w| cos(θ_z − θ_w)"